OpenStudy (anonymous):

Write the equation of a line in point-slope form that has a slope of two-thirds and passes through the point (2, -7).

OpenStudy (anonymous):

y-y1=m(x-x1) y-7=2/3(x-2)

OpenStudy (anonymous):

that should be

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

y-(-7)

OpenStudy (anonymous):

y = (2 x)/3-25/3 and -2 x+3 y+25 = 0